Mr. Levy borrowed money from two lenders. One lender charged 5.5 percent simple interest and the other 3.5 percent simple interest. Mr. Levy had to borrow $200 more at 5.5 percent than he did at 3.5 percent. If the total interest after the first year was $245, how much did he borrow at each rate?
Answer by nikolatesla1(6)

the equation would be: x+200(1+0.055)+ x(1+0.035)=245. that can be simplified to x+211+1.035x=245; combine like terms 2.035x+211=245. subtract 211 from both sides: 2.035x=34. divide both sides by 2.035 to get x by itself. so x is approximately 16.71. you add 200 since he borrowed 200 more at 5.5 interest to get that he borrowed 16.71 at 3.5 percent and borrowed 216.71 at 5.5 percent.